Find the domain and range of the function. (Enter your answers using interval notation.)
f(x) = 2x, −5 ≤ x ≤ 7

Answer by ikleyn(52798)

.

The domain is this GIVEN interval  -5 <= x <= 7  from -5 to 7, inclusive.


In the interval form, it is the set [-5,7].


The function f(x) = 2x is a linear function, monotonically increasing.


Therefore, its range is the interval between f(-5) = 2*(-5) = -10 and f(7) = 2*7 = 14.


ANSWER.  The domain is the interval [-5,7].  The range is the interval [-10,14].

Solved, with explanations.

Is it clear to you, from the first word to the last one ?